chlorine can be prepared by in the laboratory by the reaction of manganese dioxide with the hydrochloric acid HCl(ag) as described by the equation equation

MnO2(s) + 4 HCl(aq) --> MnCl2(aq) + 2H2O(l) + Cl2 (g)

how much MnO2(s) should be added to excess HCl(aq) to obtain 345 mL Cl2(g) at 25 C and 795 Torr?
Answer #1

Given:

P = 795.0 torr

= (795.0/760) atm

= 1.0461 atm

V = 345.0 mL

= (345.0/1000) L

= 0.345 L

T = 25.0 oC

= (25.0+273) K

= 298 K

find number of moles using:

P * V = n*R*T

1.0461 atm * 0.345 L = n * 0.08206 atm.L/mol.K * 298 K

n = 1.476*10^-2 mol

From reaction,

Mol of MnO2 reacted = mol of Cl2 produced

= 1.476*10^-2 mol

Molar mass of MnO2,

MM = 1*MM(Mn) + 2*MM(O)

= 1*54.94 + 2*16.0

= 86.94 g/mol

use:

mass of MnO2,

m = number of mol * molar mass

= 1.476*10^-2 mol * 86.94 g/mol

= 1.283 g

Answer: 1.28 g